A function f:[a,b]–R is defined as
f(x) = 4 if x is rational
f(x) = -3 if x is irrational
Check whether f is integrable on [a,b]. Justify your answer
1
Expert's answer
2020-12-28T18:52:49-0500

f is not Integrable if "a \neq b"


Proof :-

Let a partition be, "P"

Where "P : a=x_0<x_1<...<x_n=b"

Let, "I_j= [x_{j-1},x_j]" , "M_j= \sup_{I_j} f" , "m_j=\inf_{I_j}f"

Now "\forall j\in \{1,2,...,n\}" "\exists" a rational and an irrational number "\in I_j"

"\Rightarrow M_j=4" and "m_j=-3" "\forall j\in \{1,2,...,n\}"


"\therefore U(f, P) = \sum_{j=1}^{\infty}|I_j|M_j=\sum_{j=1}^{\infty}|I_j|\times4"

"=4\times\sum_{j=1}^{\infty}|I_j|=4(b-a)"


and "L(f, P) = \sum_{j=1}^{\infty}|I_j|m_j=\sum_{j=1}^{\infty}|I_j|\times(-3)"


"=(-3)\times\sum_{j=1}^{\infty}|I_j|=3(a-b)"



So, Lower Integral, "L_1= \sup_{P}L(f,P)=\sup_{P}3(a-b)=3(a-b)"

and, Upper Integral, "L_2= \sup_{P}U(f,P)=\inf_{P}4(b-a)=4(b-a)"


Now if "a\neq b,"

"L_1=3(a-b)\neq 4(b-a)=L_2"


Hence, "f" is not integrable on "[a,b]" , if "a\neq b"


If "a=b," clearly f is integrable and "\int_a^bf=0" .